# An Efficient Algorithm for Generating Minimal Unique-Cause MC/DC Test cases for Singular Boolean Expressions

### v3

@@ 5. Discussion
The results have three main implications: optimality and correctness, efficiency in computational complexity, and deterministic reliability. Robin’s Rule avoids the exponential cost of truth-table or exhaustive search methods by directly constructing the minimal set. Its deterministic nature is a key advantage for tool qualification where predictability and reproducibility are paramount. The primary threat to validity is the study’s focus on SBEs. While SBEs are highly prevalent (Chilenski2001, ), extending this work to Non-SBEs is necessary. Another limitation is the comparison against a single commercial tool. VectorCAST was chosen as a certified oracle to primarily validate correctness, and a broader comparative study against other tools remains an area for future work.
@@ 6. Conclusion and Future Work
This paper presented Robin’s Rule, a deterministic algorithm that directly constructs a minimal $N+1$ test set for SBEs to satisfy Unique-Cause MC/DC. Experiments on a TCAS-II based benchmark, validated with a certified tool, proved that our method consistently achieves 100% coverage with the theoretical minimum number of test cases. This work offers a practical and optimal solution for developers in safety-critical domains, enabling rigorous software verification at a reduced cost. Future work will focus on extending the algorithm to handle Non-SBEs (coupled conditions) and adapting it for the Masking MC/DC environment.

### v4

@@ 4.6. Discussion
The results have three main implications: optimality and correctness, efficiency in computational complexity, and deterministic reliability. Robin’s Rule avoids the exponential cost of truth-table or exhaustive search methods by directly constructing the minimal set, and runs in O($N^{2}$) time by directly constructing and filling the (N+1)(N+1)×NN test table. Its deterministic nature is a key advantage for tool qualification where predictability and reproducibility are paramount. The primary threat to validity is the study’s focus on SBEs. While SBEs are highly prevalent (Chilenski2001, ), extending this work to Non-SBEs is necessary. Another limitation is the comparison against a single commercial tool. VectorCAST was chosen as our coverage oracle to validate correctness, and a broader comparative study against other tools remains an area for future work.
@@ 5. Conclusion and Future Work
This paper presented Robin’s Rule, a deterministic algorithm that directly constructs a minimal $N+1$ test set for SBEs to satisfy Unique-Cause MC/DC. Experiments on a TCAS-II-based benchmark, validated using VectorCAST as an independent coverage oracle, demonstrated that our method consistently achieves 100% coverage with the theoretical minimum number of test cases. This work offers a practical and optimal solution for developers in safety-critical domains, enabling rigorous software verification at a reduced cost. Future work will focus on extending the algorithm to handle Non-SBEs (coupled conditions) and adapting it for the Masking MC/DC environment.
Appendix A Full SBE Benchmark Expressions
Listing 1: SBE No. 1 (TCAS-II-derived)
!(a && b) && (c && !d && !e $|$ !f && g && !h $|$ !i && !j && !k) && (l && m && (n $|$ o) && p $|$ q && (r $|$ s) && !t $|$ u && (v $|$ w))
Listing 2: SBE No. 2 (TCAS-II-derived)
a && (!b $|$ !c) && d $|$ e
Listing 3: SBE No. 3 (TCAS-II-derived)
a && (!b $|$ !c $|$ d && e && !(!f && g && h && !i $|$ !j && k && l) && !(!m && n && o && p $|$ !q && !r && s)) $|$ t
Listing 4: SBE No. 4 (TCAS-II-derived)
(!a && b $|$ c && !d) && !(e && f) && !(g && h) && !(i && j) && ((k && l $|$ m && n) && o && (!p $|$ !q && !r $|$ !s && (!t $|$ !u)))
Listing 5: SBE No. 5 (TCAS-II-derived)
(!a && b $|$ c && !d) && !(e && f) && !(g && h) && ((i && j $|$ k && l) && m && (n && o $|$ !p && q))
Listing 6: SBE No. 6 (TCAS-II-derived)
!(a && b) && (!c && d && !e && !f && (g && h $|$ !i && j))
Listing 7: SBE No. 7 (TCAS-II-derived)
a && !b && !c && d && !e && f && (g $|$ !h && (i $|$ j)) && !(k && l $|$ !m && n $|$ o)
Listing 8: SBE No. 8 (TCAS-II-derived)
a && !b && !c && !((d && (e $|$ !f && (g $|$ h))) $|$ i && (j $|$ !k && (l $|$ m)) && !n && !o) && !(p && q $|$ !r && s && !t)
Listing 9: SBE No. 9 (TCAS-II-derived)
a && !b && !c && (d && (e $|$ !f && (g $|$ h)) && (!i && !j $|$ k) $|$ !l) && (m && n $|$ !o && p && !q)
Listing 10: SBE No. 10 (TCAS-II-derived)
a $|$ b $|$ c $|$ !d && !e && f && g && !h && !i $|$ j && (k $|$ l) && !m
Listing 11: SBE No. 11 (TCAS-II-derived)
a && b && (c $|$ d) && e $|$ f && (g $|$ h) && !i $|$ j && (k $|$ l)
Listing 12: SBE No. 12 (TCAS-II-derived)
a && ((b $|$ c $|$ d) && e $|$ f && g $|$ h && (i $|$ j $|$ k $|$ l)) $|$ (m $|$ n) && (o $|$ p $|$ q) && r
Listing 13: SBE No. 13 (TCAS-II-derived)
(a && b $|$ c && d) && e && (f $|$ (g && (h && i $|$ j && k)))
Listing 14: SBE No. 14 (TCAS-II-derived)
(a && b $|$ c && d) && e && (f && g $|$ !h && i)
Listing 15: SBE No. 15 (TCAS-II-derived)
!a && b && !c && !d && (e && f $|$ !g && h)
Listing 16: SBE No. 16 (Randomly generated)
(((a $|$ b) $|$ c && !(!(d && e)) && f $|$ g) $|$ !h $|$ i $|$ !((j $|$ !(!((k && (l && !m))) $|$ n && o $|$ p $|$ q && r $|$ s))))
Listing 17: SBE No. 17 (Randomly generated)
((a && b && c && d) && !(e $|$ f) $|$ (!(!g) && !h && i) && j && k)
Listing 18: SBE No. 18 (Randomly generated)
!(a $|$ b) $|$ (c && d && !e $|$ !f && g $|$ (!h $|$ i) && j && !k) $|$ (l && m && (n $|$ o) && p $|$ q && (r $|$ s) && !t && u && (v && w) $|$ x)
Listing 19: SBE No. 19 (Randomly generated)
!(a $|$ (b && !c)) && (!(!(d && e)) && f $|$ g $|$ (!((h $|$ i) $|$ !j) && !(!k))) $|$ !(!(l && m)) && n && o && !p $|$ (!q $|$ r)
Listing 20: SBE No. 20 (Randomly generated)
((!((a && b)) && !((c && (!(!d) $|$ e && f)) && ((!g && (!h $|$ !(!i) $|$ (j $|$ k))) && !l $|$ m) && n $|$ !(!o) $|$ p)) $|$ !(!((q $|$ r $|$ s) $|$ !t && !((!(!u) $|$ v $|$ w)) $|$ !(!x) && !y && z)))
Listing 21: SBE No. 21 (Randomly generated)
(a $|$ !b $|$ !(c && d) $|$ (e && f) $|$ g) && (h && i && j) $|$ (k && l && m)
Listing 22: SBE No. 22 (Randomly generated)
a $|$ b && c $|$ d && e $|$ (!(!(f $|$ g && h) $|$ i && !(!j) && !k) && l $|$ m $|$ !n && o && p && q)
Listing 23: SBE No. 23 (Randomly generated)
(!a && b $|$ c) $|$ d && e $|$ !f && g
Listing 24: SBE No. 24 (Randomly generated)
a && !b && c $|$ d && !e && !((f && g $|$ (!h && i)) && j $|$ k)
Listing 25: SBE No. 25 (Randomly generated)
!((a && b)) && c $|$ (!(((d $|$ e) $|$ f) && g) && h && !(!(i && j))) $|$ k && l $|$ m $|$ n $|$ o && (p && q && r)